OpenStudy (anonymous):

Deandre races his bicycle for 124m . A wheel of his bicycle turns 31 times as the bicycle travels this distance. What is the diameter of the wheel? Use the value 3.14 for Pi . Round your answer to the nearest tenth. Do not round any intermediate steps. Deandre races his bicycle for 124m . A wheel of his bicycle turns 31 times as the bicycle travels this distance. What is the diameter of the wheel? Use the value 3.14 for Pi . Round your answer to the nearest tenth. Do not round any intermediate steps. @Mathematics

OpenStudy (saifoo.khan):

circumfrence = 2 p i r

OpenStudy (anonymous):

im confused whats the answer

OpenStudy (saifoo.khan):

0.6m

OpenStudy (saifoo.khan):

124/ 31 = 4 circumfrence = 4 4 = 2 pi r 2 = pr r r = 2/pi r = 0.6m